OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 3x

OpenStudy (akashdeepdeb):

sin3x can be written as sin 3x = sin(2x+x) Now use the identity for sin(A+B) = sinA.cosB + sinB.cosA sin(2x+x) = sin2x.cosx + cos2x.sinx = 2sinx.cosx.cosx + (cos^2 x - sin^2 x)sinx Now simplify the above using identities!! :D Understood? :)

OpenStudy (anonymous):

2 sin x cos^2x + cos x 2 sin x cos^2x + sin^3x sin x cos^2x - sin^3x + cos^3x 2 cos^2x sin x + sin x - 2 sin^3x

OpenStudy (anonymous):

those are my options